OpenStudy (anonymous):

Use cross products to determine which pair of ratios forms a proportion. a.2/3=10/18 b.2/3=12/18 c.2/3=14/15 d.2/3=14/24

OpenStudy (anonymous):

we are going to have to cross multiply...do you know how to do that ?

OpenStudy (anonymous):

hello ??

OpenStudy (anonymous):

yeah.sorry i was trying to figure it out lol.but its hard.yes i know how to cross multiply.

OpenStudy (anonymous):

I will do the first one for you.... (a) 2/3 = 10/18 (cross multiply) (2)(18) = (3)(10) 36 = 30 (incorrect) this is not proportional because they do not equal.

OpenStudy (anonymous):

ohkay so.(2)(18)=(3)(12) 36=36.so this is the answer!(:

OpenStudy (anonymous):

yes....you got it. Answer B

OpenStudy (anonymous):

good job :)